OpenStudy (anonymous):

(7/x) - (1/7) = (2/x)

OpenStudy (zzr0ck3r):

I would multiply every term by x and see what you get

OpenStudy (lgbasallote):

\[\frac 7x - \frac 17 = \frac 2x\] you can start by adding \(\frac 17\) to both sides \[\implies \frac 7x = \frac 2x + \frac 17\] now subtract \(\frac 2x\) from both sides \[\implies \frac 7x - \frac 2x = \frac 17\] if you notice.. \(\frac 7x\) and \(\frac 2x\) have common denominators so i can just subtract the numerators then vopy the denominator \[\implies \frac{7-2}{x} = \frac 17\] combine like terms \[\implies \frac 5x = \frac 17\] now cross multiply \[\implies 5 \times 7 = x\] does that help?

OpenStudy (zzr0ck3r):

jinks

OpenStudy (lgbasallote):

lol i had a different method

OpenStudy (anonymous):

Thanks igbasallote. I have a few more questions. Can you help me?

OpenStudy (lgbasallote):

only if you understood what i wrote above

OpenStudy (anonymous):

I do for sure. I just solved one just like that but I had to factor instead. I just need help with reviewing Algebra II Trig. I can't remember a thing

OpenStudy (lgbasallote):

so what's your question?

OpenStudy (anonymous):

Find an equation of the vertical line through the point (-7.7, 1.9). Can you give me the equation and then I could see if I can figure it out myself?

OpenStudy (lgbasallote):

here's a hint: the equation of a vertical line is x = something

OpenStudy (anonymous):

X = 0?

OpenStudy (lgbasallote):

here's another hint... the equation of the HORIZONTAL line that passes through (1,2) is y = 2 the equation of the VERTICAL line that passes through (1,2) is x = 1

OpenStudy (anonymous):

So it's just x = -7.7?

OpenStudy (lgbasallote):

yup. you learn fast

OpenStudy (anonymous):

That was just a dumb question on my part. Summer made me forget everything. I have a few more questions and I promise I will stop bugging you

OpenStudy (lgbasallote):

lol. sure ask away

OpenStudy (anonymous):

Find an equation for the line with the given properties. Express the answer using the slope-intercept form of the equation of the line. X-intercept = 3; y-intercept = -4 Please keep giving me hints so I learn sorta on my own

OpenStudy (lgbasallote):

if the x-intercept is 3...do you know how it looks like in ordered pair form?

OpenStudy (anonymous):

What do you mean? (3,-4)?

OpenStudy (lgbasallote):

when you see something like x-intercept = 1 this means the point is (1,0) when you see something like y-intercept = 2 this means the point is (0,2) does that help?

OpenStudy (anonymous):

Yea I understand that. But what does that have to do with finding the question? I don't have the slope or anything.

OpenStudy (lgbasallote):

when you are given two points and you are asked to find the equation of the line, you just use this formula \[\Large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] are you familiar with those notations?

OpenStudy (anonymous):

Yea ok I couldn't remember that equation. Lgive me a second to see if I can solvent and get it right

OpenStudy (anonymous):

I can't remember. I have 4= (4/3) * 3

OpenStudy (lgbasallote):

y and x stay...they are variables..you dont substitute anything to them

OpenStudy (lgbasallote):

you just substitute x1, x2, y1 and y2

OpenStudy (anonymous):

Y=4/3x?

OpenStudy (lgbasallote):

that wasnt what i meant....

OpenStudy (lgbasallote):

first tell me what is x1?

OpenStudy (anonymous):

X1=3

OpenStudy (lgbasallote):

right. and what is x2?

OpenStudy (anonymous):

0?

OpenStudy (lgbasallote):

yes. y1?

OpenStudy (anonymous):

-4,0

OpenStudy (lgbasallote):

nope..

OpenStudy (lgbasallote):

if the y-intercept is -4 then the ordered pair is (0,-4) y1 is 0 and y2 is -4 got it?

OpenStudy (anonymous):

Yea that's what I meant. Sorry. So what is the slope?

OpenStudy (lgbasallote):

remember i said the formula is \[y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] so if you substitute those you'll get \[y - 0 = \frac{-4 - 0}{0 - 3} (x- 3)\] got it?

OpenStudy (anonymous):

So it's y = 4/3x + 4?

OpenStudy (anonymous):

I mean -4

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

Ok thanks. Next questionn I just want to check to see if I got it right.

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

Find the intercepts of the graph of the equation. 1/6x + 1/3y = 1 My equation is y = -1/2x + 3Are the intercepts (0,3)?

OpenStudy (lgbasallote):

wait... what's 1/6 x + 1/3 y = 1?

OpenStudy (lgbasallote):

that's the original equation?

OpenStudy (anonymous):

Yep

OpenStudy (lgbasallote):

hmm this is probably the best time for me to teach you another equation...the intercept formula.. when you have something like \[\frac xa + \frac yb = 1\] the intercepts are just (a,0) and (0,b) got it?

OpenStudy (anonymous):

So my fractions 1/6 and 1/3 are x and y?

OpenStudy (lgbasallote):

no.. 1/6 x is x/6 => x/a 1/3 y is y/3 => y/b

OpenStudy (anonymous):

Ok so intercepts are (6,0) and (0,3) it matches my equation I said earlier when you graph them.

OpenStudy (lgbasallote):

yup that is right

OpenStudy (lgbasallote):

actually you can also use the intercept formula on that problem you had earlier about the equation with x-intercept 3 and y-intercept -4

OpenStudy (lgbasallote):

so are we done here @Brybro ?

OpenStudy (anonymous):

Perpendicular to the line x - 7y = 7 That equation is simply y = 1/7x + 1 in simplistic form right? So the slope would be 7?

OpenStudy (lgbasallote):

close... the perpendicular slope would be NEGATIVE reciprocal

OpenStudy (anonymous):

Graph the circle X^2 + (y-1)^2 = 36 I have no clue what to do here

OpenStudy (lgbasallote):

first find the center and radius of that equation

OpenStudy (anonymous):

How do you do that?

OpenStudy (lgbasallote):

the equation of a circle is [(x-h)^2 + (y-k)^2 = r^2\] (h,k) is the center and r is the radius

OpenStudy (lgbasallote):

equation of a circle is \[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

So 1 is the center and 1 is the radius?

OpenStudy (lgbasallote):

?

OpenStudy (lgbasallote):

i said (h,k) is the center and r is the radius...

OpenStudy (anonymous):

Oh. My bad. So center is (1,1) and radius is 6?